A NON ABELIAN TOOL FOR LOCAL-TO-GLOBAL PROBLEMS

RONALD BROWN

A 1-fold groupoid is just a groupoid and and n-fold groupoid is a set with n groupoid structures any two of which satisfy an interchange law.

The aim of the talk is to show the intuitions behind how how types of n-fold groupoid have been used in homotopy theory and the cohomology of groups to give new calculations, with a view to encouraging their use in differential situations.